A parabola has a directris at x=3 and a focus at (-4,3). Find its equation in standard form , and please provide its x-intercept. (:
Answer by lwsshak3(11628)

A parabola has a directrix at x=3 and a focus at (-4,3). Find its equation in standard form , and please provide its x-intercept.
..
Standard form of parabola:(y-k)^2=4p(x-h), with (h,k) being the (x,y) coordinates of the vertex.
For given parabola:
vertex at (-1/2,3)
axis of symmetry y=3
p=1/2
Equation: (y-3)^2=-2(x+1/2)
x-intercept
set y=0
(-3)^2=-2x-1
9=-2x-1
-2x=10
x=-5 (x-intercept)